# Distributed Deep Reinforcement Learning Based Gradient Quantization for Federated Learning Enabled Vehicle Edge Computing

### v1

@@ I Introduction
With the rapid development of autonomous driving technology, a large amount of data are generated by various sensors on vehicles, such as cameras, radar, lidar, as well as proximity and temperature sensors[1]. For example, a self-driving car is expected to generate about 1 GB of data per second[2]. Vehicles need powerful computing capability to process and analyze the data to support the modeling training of the vehicular artificial intelligence (AI) applications such as simultaneous localization and mapping (SLAM), augmented reality (AR) navigation, object tracking, and high-definition (HD) map generation[3]. However, the computing capability of vehicles is limited. In this situation, vehicle edge computing (VEC) becomes a promising technology to facilitate these applications, where a base station (BS) connected with an edge server can collect and utilize the vehicles’ data for model training[4, 5, 6]. However, the raw data generated by a vehicle often contains personal information, thus there may be a risk of data leakage in data privacy in VEC[7].
@@
Recently, federated learning (FL) has been enabled in VEC[8]. FL can protect the privacy of vehicles to a certain extent through sharing the gradients of vehicles’ local models instead of local data[9]. FL training process consists of multiple rounds. For each round, vehicles individually train their local models using local data, then the BS collects the gradients of local models and aggregates them to obtain a global model, after that the edge server sends the global model to vehicles for the next round training. After meeting the convergence requirement, i.e., reaching the desired precision or the predefined loss value, the entire training process is completed[10]. The gradients of vehicles’ local models are usually large for the vehicular AI applications, thus transmitting such large gradients would cause large per-round latency[11, 12].
@@
Gradient quantization may reduce the per-round latency in FL enabled VEC through compressing gradients and reducing the number of bits, i.e., the quantization level, to transmit gradients. In gradient quantization, how to determine the quantization level is a key issue[13]. On one hand, a few quantization level may significantly reduce the transmission overhead, thus cause a low per-round latency and further result in a small total training time. On the other hand, lower number of quantization levels may produce bigger quantization error (QE) due to the heavily compression, which may increase the number of rounds required for convergence and increase the total training time[14, 15]. That means there exists a reasonable gradient quantization level which can adjust the trade off between the total training time and QE for the FL enabled VEC. If only the total training time is optimized, it may lead to the excessive pursuit of large quantization error and the sacrifice of model performance in order to reduce the number of communication rounds, and it is difficult to guarantee the convergence quality and accuracy of the model[16]. Therefore, it is critical to jointly optimize the total training time and QE for the FL enabled VEC. To best of our knowledge, there is no work presented, which motivates us to conduct this work.
@@
In the FL enabled VEC, the channel condition is time-varying due to the time-varying path loss caused by the mobility of the vehicle, which causes challenges to conduct this work[17, 18]. Deep reinforcement learning (DRL) is a favorable framework to solve such problem in complex environments[19]. There are existing papers adopting the centralized DRL framework to dynamically adjust the quantization level in wireless networks[20, 21]. However, the centralized DRL framework lead to significantly high traffic load. In addition, for the centralized DRL framework, the high mobility of vehicles makes the BS difficult to collect accurate instantaneous channel state information (CSI)[22]. On the contrary, in the distributed DRL framework, each vehicle makes decisions independently and does not need to send all the information to the central node, which can reduce the communication overhead and can reduce the network traffic blocking to the BS at the same time[23]. Since each vehicle makes decisions based on its own state, distributed DRL can better collect the CSI[24, 25]. Based on the above analysis, it is necessary to find a proper distributed DRL framework to allocate the quantization levels in the FL enabled VEC.
@@
In this paper, we consider the time-varying channel conditions and propose a distributed DRL-based quantization level allocation scheme by optimizing the long-term reward in terms of the total training time and QE [^1] The source code has been released at: https://github.com/qiongwu86/Distributed Deep Reinforcement Learning Based Gradient Quantization for Federated Learning Enabled Vehicle Edge Computing. The main contributions of this paper are summarized as follows.
@@ III System Model
We consider a quantized FL system consisting of $N$ vehicles and a BS connected with an edge server. The vehicles are driving in the coverage of the BS. The velocities of different vehicles follow Gaussian distribution and the velocity of each vehicle is not changed. The driving duration of the vehicles is divided into time slots[44]. The set of the vehicles is denoted by $\mathcal{V}=\{V_{1},V_{2},V_{3},\vbox{\hbox{$\cdots$}},V_{n},\vbox{\hbox{$\cdots$}},V_{N}\}$. Each vehicle $V_{n}$ is equipped with a MEC processor that provides the local processing capability. Vehicles get the data samples through sensors or communication, and each vehicle $V_{n}$ holds the local dataset $\mathcal{D}_{n}$ with a uniform size of $D$, the dataset $\mathcal{D}_{n}$ contains data samples $\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{D}$, where $\mathbf{x}_{i}$ represents the feature of the data sample and $y_{i}$ represents the label of the data sample[45]. The features of the data sample are used as the input of the model training, and the labels of the data sample are the expected outputs corresponding to the input data, which are used to train and validate the model[46, 47, 48].
@@ 23            The algorithm stops
Each vehicle $V_{n}$ calculates the utility $\varphi_{n}^{r}$ to determine if vehicle $V_{n}$ is selected to participate in round $r$. Vehicles with similar local models to the global model contribute less to the training of the current global model and do not need to participate in the training frequently. Meanwhile, the selected vehicles need to have enough time to participate in the current round of training and upload the local model[51]. Thus, $\varphi_{n}^{r}$ is affected by both model similarity $\alpha_{n}^{r}$ and vehicle’s residence time in the coverage $\beta_{n}^{r}$, i.e., (line 9)
@@ V-C Optimization problem
The optimization problem P1 is challenging to solve using conventional methods because it is a non-linear constrained integer optimization problem[58]. Although some techniques, such as the branch and bound method, can provide approximate solutions, their computational complexities are very high[59]. Moreover, the conventional optimization methods are based on the assumption that the channel state information (CSI) is known. However, it is impractical to obtain all CSI information in the real environment[60]. To address these challenges, DRL can be employed to handle imperfect information. By interacting with dynamic channel environments, DRL can train a model by learning from online user behavior and environmental conditions, and thus maximizes the system utility[61, 62]. This makes DRL a promising solution for tackling these problems. Since the centralized DRL framework results in significantly high traffic load, and the high mobility of vehicles makes it difficult for the BS to collect accurate CSI. In the next section, we will propose a distributed DRL scheme to solve the optimization problem P1.

### v2

@@ I Introduction
With the rapid development of autonomous driving technology, a large amount of data are generated by various sensors on vehicles, such as cameras, radar, lidar, as well as proximity and temperature sensors. For example, a self-driving car is expected to generate about 1 GB of data per second[1]. Vehicles need powerful computing capability to process and analyze the data to support the modeling training of the vehicular artificial intelligence (AI) applications such as simultaneous localization and mapping (SLAM), augmented reality (AR) navigation, object tracking, and high-definition (HD) map generation[2]. However, the computing capability of vehicles is limited. In this situation, vehicle edge computing (VEC) becomes a promising technology to facilitate these applications, where a base station (BS) connected with an edge server can collect and utilize the vehicles’ data for model training[3]. However, the raw data generated by a vehicle often contains personal information, thus there may be a risk of data leakage in data privacy in VEC[4].
@@
Recently, federated learning (FL) has been enabled in VEC[5]. FL can protect the privacy of vehicles to a certain extent through sharing the gradients of vehicles’ local models instead of local data[6]. FL training process consists of multiple rounds. For each round, vehicles individually train their local models using local data, then the BS collects the gradients of local models and aggregates them to obtain a global model, after that the edge server sends the global model to vehicles for the next round training. After meeting the convergence requirement, i.e., reaching the desired precision or the predefined loss value, the entire training process is completed[7]. The gradients of vehicles’ local models are usually large for the vehicular AI applications, thus transmitting such large gradients would cause large per-round latency.
@@
Gradient quantization may reduce the per-round latency in FL enabled VEC through compressing gradients and reducing the number of bits, i.e., the quantization level, to transmit gradients. In gradient quantization, how to determine the quantization level is a key issue. On one hand, a few quantization level may significantly reduce the transmission overhead, thus cause a low per-round latency and further result in a small total training time. On the other hand, lower number of quantization levels may produce bigger quantization error (QE) due to the heavily compression, which may increase the number of rounds required for convergence and increase the total training time[8, 9]. That means there exists a reasonable gradient quantization level which can adjust the trade off between the total training time and QE for the FL enabled VEC. If only the total training time is optimized, it may lead to the excessive pursuit of large quantization error and the sacrifice of model performance in order to reduce the number of communication rounds, and it is difficult to guarantee the convergence quality and accuracy of the model. Therefore, it is critical to jointly optimize the total training time and QE for the FL enabled VEC. To best of our knowledge, there is no work presented, which motivates us to conduct this work.
@@
In the FL enabled VEC, the channel condition is time-varying due to the time-varying path loss caused by the mobility of the vehicle, which causes challenges to conduct this work[10]. Deep reinforcement learning (DRL) is a favorable framework to solve such problem in complex environments[11]. There are existing papers adopting the centralized DRL framework to dynamically adjust the quantization level in wireless networks[12, 13]. However, the centralized DRL framework lead to significantly high traffic load. In addition, for the centralized DRL framework, the high mobility of vehicles makes the BS difficult to collect accurate instantaneous channel state information (CSI). On the contrary, in the distributed DRL framework, each vehicle makes decisions independently and does not need to send all the information to the central node, which can reduce the communication overhead and can reduce the network traffic blocking to the BS at the same time. Since each vehicle makes decisions based on its own state, distributed DRL can better collect the CSI[14, 15]. Based on the above analysis, it is necessary to find a proper distributed DRL framework to allocate the quantization levels in the FL enabled VEC.
@@
In this paper, we consider the time-varying channel conditions and propose a distributed DRL-based quantization level allocation scheme by optimizing the long-term reward in terms of the total training time and QE [^1] The source code has been released at: https://github.com/qiongwu86/Distributed-DRL-Based-Gradient-Quantization-for-Federated-Learning-Enabled-VEC. The main contributions of this paper are summarized as follows.
@@ III System Model
We consider a quantized FL system consisting of $N$ vehicles and a BS connected with an edge server. The vehicles are driving in the coverage of the BS. The velocities of different vehicles follow Gaussian distribution and the velocity of each vehicle is not changed. The driving duration of the vehicles is divided into time slots[34]. The set of the vehicles is denoted by $\mathcal{V}=\{V_{1},V_{2},V_{3},\vbox{\hbox{$\cdots$}},V_{n},\vbox{\hbox{$\cdots$}},V_{N}\}$. Each vehicle $V_{n}$ is equipped with a MEC processor that provides the local processing capability. Vehicles get the data samples through sensors or communication, and each vehicle $V_{n}$ holds the local dataset $\mathcal{D}_{n}$ with a uniform size of $D$, the dataset $\mathcal{D}_{n}$ contains data samples $\left\{(\mathbf{x}_{i},y_{i})\right\}_{i=1}^{D}$, where $\mathbf{x}_{i}$ represents the feature of the data sample and $y_{i}$ represents the label of the data sample. The features of the data sample are used as the input of the model training, and the labels of the data sample are the expected outputs corresponding to the input data, which are used to train and validate the model[35, 36].
@@ 23            The algorithm stops
Each vehicle $V_{n}$ calculates the utility $\varphi_{n}^{r}$ to determine if vehicle $V_{n}$ is selected to participate in round $r$. Vehicles with similar local models to the global model contribute less to the training of the current global model and do not need to participate in the training frequently. Meanwhile, the selected vehicles need to have enough time to participate in the current round of training and upload the local model. Thus, $\varphi_{n}^{r}$ is affected by both model similarity $\alpha_{n}^{r}$ and vehicle’s residence time in the coverage $\beta_{n}^{r}$, i.e., (line 9)
@@ V-C Optimization problem
The optimization problem P1 is challenging to solve using conventional methods because it is a non-linear constrained integer optimization problem. Although some techniques, such as the branch and bound method, can provide approximate solutions, their computational complexities are very high. Moreover, the conventional optimization methods are based on the assumption that the channel state information (CSI) is known. However, it is impractical to obtain all CSI information in the real environment[45]. To address these challenges, DRL can be employed to handle imperfect information. By interacting with dynamic channel environments, DRL can train a model by learning from online user behavior and environmental conditions, and thus maximizes the system utility[46]. This makes DRL a promising solution for tackling these problems. Since the centralized DRL framework results in significantly high traffic load, and the high mobility of vehicles makes it difficult for the BS to collect accurate CSI. In the next section, we will propose a distributed DRL scheme to solve the optimization problem P1.